OpenStudy (anonymous):

Use the quadratic formula to find the zeros of y=2x^2+8x+1

Parth (parthkohli):

\[x = {-b \pm \sqrt{b^2 - 4ac} \over 2a}\]

OpenStudy (anonymous):

thats not one of the options @ParthKohli

OpenStudy (anonymous):

@Coolsector

OpenStudy (anonymous):

@yahoo!

OpenStudy (anonymous):

you should use what parthkohli wrote .. ax^2+bx+c =0 use what he wrote

OpenStudy (anonymous):

it's not an option though @Coolsector

OpenStudy (anonymous):

no .. look ax^2+bx+c =0 in your case a=2 b =8 c=1 now plug it into what parthkoli wrote and you will get the zeros ..

OpenStudy (anonymous):

Look at this.

OpenStudy (anonymous):

@Coolsector

OpenStudy (anonymous):

if you plug it into what partkholi wrote you get : \[\frac{ -8 \pm \sqrt{64-8} }{ 4 }\]